find three consecutive odd integers such that the product of the second and third is 55 less than 30 times the first
Answer by lwsshak3(11628)

find three consecutive odd integers such that the product of the second and third is 55 less than 30 times the first
***
let x=1st consecutive odd integer
x+2=2nd consecutive odd integer
x+4=3rd consecutive odd integer
...
(x+2)(x+4)=30x-55
x^2+6x+8=30x-55
x^2-24x+63=0
(x-21)(x-3)=0
x=21
x+2=23
x+4=25
or
x=3
x+2=5
x+4=7
2 sets of 3 consecutive odd integers:
21,23,25
and
3,5,7